# Self-Calibrated CLIP for Training-Free Open-Vocabulary Segmentation

### v2

@@ 3.3 Self-Adjusting for Semantic Coherence
After resolving the anomaly tokens, the model reduces focus on them in the last layer’s attention. However, a challenge still remains: anomaly tokens have already caused substantial disruption to the other normal tokens in the previous layers, diminishing their local awareness.
@@
To enhance feature discriminability and correlation, some methods [26, 54] leverage backbones with strong spatial coherence like DINO [7], to provide fine-grained details. Although this approach improves performance, it introduces an additional backbone and incurs extra computational costs during either the training or testing phase.
@@
Figure 4: Top: Visualization of patch similarities shows CLIP’s deep layers perform poorly, but its mid layers exhibit semantic consistency comparable to DINO. Bottom Left: ROC curve analysis further supports this, with SC-CLIP showing superior semantic coherence. Bottom Right: Detailed ROC analysis of our method.
@@
Figure 5: Illustration of the self-adjusting strategy. (a) We use the similarity map from CLIP’s mid layer to adaptively aggregate deep features by combining semantically similar patches, resulting in a clearer segmentation map. The second row provides a detailed view of the process for the selected patch $\bigstar$. (b) We apply the similarity map to enhance attention, broadening and refining the activation regions.
@@
This motivates us to explore whether such semantic coherence can be uncovered within CLIP itself. We begin by visualizing patch similarities, as shown in Figure 4 (top), where we observe the CLIP’s last layer performs poorly as expected. However, its mid-layer exhibits stronger semantic coherence, comparable to that of DINO. To further validate this observation, we quantitatively evaluate all features of CLIP using ROC curve analysis. Specifically, using the ADE20K dataset [73], we extract patches from each layer $l$ and compute cosine similarity between them to create the similarity map $\mathbf{Simi}^{l}\in\mathbb{R}^{N\times N}$, defined as $\mathbf{Simi}^{l}=\frac{\mathbb{X}^{l}\cdot\mathbb{X}^{l}}{\|\mathbb{X}^{l}\|\|\mathbb{X}^{l}\|}$. This similarity map serves as a binary classifier to indicate whether two patches belong to the same category, with patches of the same category labeled as 1 and otherwise 0. Each patch’s category is determined by majority voting on pixel labels in the segmentation map. For comparison, we also evaluate DINO. A higher area under the curve (AUC) in ROC analysis reflects better semantic consistency. As shown in Figure 4 (bottom left), CLIP’s mid-layer features exhibit strong spatial coherence (AUC=0.76), closely matching DINO (AUC=0.77). However, CLIP’s last-layer features perform much worse (AUC=0.66).
@@
The aggregated features $\hat{\mathbb{X}}^{deep}$ show a clearer and more precise segmentation results compared to the original $\mathbb{X}^{deep}$.
@@
Moreover, we find that the self-self attention (e.g., K-K attention) proposed in the previous works [25, 49, 29, 4] exhibit insufficient attention activations. Therefore, we further incorporate the similarity map $\mathbf{Simi}^{mid}$ to augment the attention operation, as detailed in the following formula. As shown in Figure 5 (b), the attention regions now display more extensive and accurate activations.
[…]
@@
Feature aggregation and attention enhancement constitute the self-adjusting strategy, which significantly improves the semantic coherence, increasing the AUC to 0.80, as shown in Figure 4 (bottom right). This demonstrates that the inherent consistency within CLIP’s intermediate features can be effectively leveraged for adjustment.
3.4 Multi-level Feature Fusion
@@
Motivated by the effectiveness of multi-level feature aggregation in enriching details for dense prediction tasks [32, 30, 57, 17, 29], we explore the potential of leveraging CLIP’s multi-level features in a training-free manner. A straightforward approach is to directly sum multi-level features: $\mathbb{X}^{last}+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$, where $\mathcal{M}$ is the set of multi-level features. However, CLIP exhibits significant discrepancies across different layers. Specifically, the similarity between $\mathbb{X}^{last}$ and $\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ is merely 0.094. Directly summing them severely disrupts CLIP’s cross-modal alignment capability. We identify two key principles for training-free feature fusion: 1) Ensuring the compatibility between $\mathbb{X}^{last}$ and $\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ is crucial. This can be achieved by leveraging the parameter space of the last layer for alignment. The similarity between $\mathbb{X}^{last}$ and $\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right)$ increases to 0.983, where $\mathbf{L}$ denotes the last layer. This shows that using the last layer can significantly enhance feature compatibility. 2) Maintaining the integrity of $\mathbb{X}^{last}$, as it contains the direct alignment with the text embeddings.
@@
Based on these insights, we propose a two-pass strategy, where CLIP’s final layer is explicitly employed for alignment. This strategy involves two forward passes—one with the original $\mathbb{X}^{penul}$ and another with multi-level features, formulated as: $\mathbf{L}(\mathbb{X}^{penul})+\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right).$ Ablation experiments validate the effectiveness of our approach.
4 Experiments
4.1 Experiment Settings
@@
Datasets and Metric We conduct comprehensive evaluations on eight commonly used benchmark datasets, which are grouped into two categories: 1) with a background class, including PASCAL VOC (VOC21) [15], PASCAL Context (Context) [38], and COCO Object (COCO-Obj) [6]; 2) without a background class, including PASCAL VOC20 (VOC20) [15], Cityscapes (City) [11], PASCAL Context59 (Context59) [38], ADE20K (ADE) [73], and COCOStuff (COCO-Stf) [6]. And we evaluate results with the standard mean-intersection-over-union (mIoU) as the metric.
@@
Table 1: Performance comparison of our approach with other methods on eight semantic segmentation benchmarks. For a fair comparison, we reproduce all methods following the evaluation protocol in Section 4.1, considering the different settings used by each method. CLIP-DINOiser denotes our reproduced results using OpenAI’s pretrained weights. We report ProxyCLIP* results using its DINO-B/16 variant.
@@
Implementation Details In our experiments, we utilize the CLIP [41] with ViT-B/16 and ViT-L/14 architectures. And our code implementation is built on the MMSegmentation [10]. For the evaluation protocol, we adopt the sliding window inference strategy from SCLIP [49]: input images are resized to have a short side of 336 (560 for Cityscapes [11] due to its higher resolution), and the slide inference is conducted with a $224\times 224$ window and a $112\times 112$ stride. No post-processing strategies are applied. Across all datasets, we use the standard ImageNet prompts [41] combined with their category names to construct text descriptions. For all the hyper-parameters, we keep consistent across all datasets without separate tuning.
4.2 Main Results
@@
To ensure a fair comparison, we strictly follow the evaluation protocol specified in the Section 4.1 to reproduce all methods. For CLIP-DINOiser [54], we retrain the model using the OpenAI’s pretrained weights. Table 1 presents the comparison of all methods, where our SC-CLIP achieves an average mIoU of 43.9% on CLIP ViT-B/16, setting a new state-of-the-art results across eight benchmarks with a notable 9.5% improvement over previous methods. Furthermore, our method demonstrates robustness across different backbones, achieving optimal performance on CLIP ViT-L/14, improving the results by 3.5% average mIoU.
@@
The vanilla CLIP achieves only a 14.4% and 6.6% mIoU on ViT-B/16 and ViT-L/14 respectively, demonstrating its limitations in capturing fine-grained spatial details. Our training-free method boosts its performance by threefold on the ViT-B/16 and by 6.8 times on the ViT-L/14.
@@
In contrast to ProxyCLIP [26], which relies on the DINO-B/16 [7] backbone to provide attention weights, our method does not depend on any auxiliary backbone yet achieves superior results. This highlights the inherent properties of CLIP can be effectively leveraged to calibrate itself, thereby enhancing its feature representation.
4.3 Ablation Study
@@
In this section, we conduct comprehensive ablation experiments to validate the effectiveness of our method. We use SCLIP [49] as our baseline, which modifies the attention mechanism in the final layer by replacing the original $\mathbf{Q}\mathbf{K}^{\top}$ attention with $\mathbf{Q}\mathbf{Q}^{\top}+\mathbf{K}\mathbf{K}^{\top}$ attention, to enhance correlations. Additionally, we remove the residual connections and feed-forward network (FFN) following [25, 26].
Table 2: Ablation experiments on the proposed strategies.
@@
+ MultiFuse 64.6 36.8 37.7 40.1 26.6 41.2
Analysis of Various Strategies
In Table 2, we incrementally incorporate each strategy to highlight its contribution. First, by resolving the anomaly tokens (denoted by AnomRes), our method achieves a 1.2% mIoU improvement, effectively addressing the issues brought by anomaly tokens, as discussed in Section 3.2. Next, we apply the self-adjusting strategy proposed in Section 3.3, which leverages the spatial consistency within CLIP’s mid-level features to enhance attention correlation (denoted by AttnEnh), leading to a 0.9% improvement. By adaptively aggregating deep features (FeatAgg), we obtain a further 0.8% gain in mIoU. Finally, the multi-level fusion strategy (MultiFuse) in Section 3.4 provides a 1.6% boost in mIoU. Collectively, these strategies contribute to a substantial 12.3% improvement over the baseline. Besides, we conduct a detailed ROC curve analysis as shown in Figure 4 (bottom right). SC-CLIP achieves an AUC of 0.81, surpassing the baseline by 17.4% and significantly enhancing the semantic coherence.
Table 3: Ablation study of resolving the anomaly tokens.
@@
Table 4: Ablation study of the attention enhancement.
@@
only $\mathbf{K}\mathbf{K}^{\top}$ 60.6 33.1 34.3 36.6 24.2 37.8
only $\mathbf{Simi}^{mid}$ 60.4 33.8 35.8 37.1 24.8 38.3
both 61.2 34.2 36.2 37.4 25.0 38.8
Resolving the Anomaly Tokens
@@
As we use the LOF algorithm to identify the anomaly tokens, we need to adjust the contamination hyperparameter to control the number of detected anomalies. We explore the optimal number of tokens to resolve, as shown in Table 3. As the number increases from 1 to 10, performance gradually improves. However, further increasing this number to 15 yields no additional gains. Therefore, we adopt the removal of 10 tokens (about 5% of the ViT-B/16 token sequence) in our method.
Self-adjusting Strategy
Table 4 shows the ablation study on attention enhancement. When both self-self attention ($\mathbf{K}\mathbf{K}^{\top}$) and $\mathbf{Simi}^{mid}$ are applied (Eq. (5)), the performance improves compared to using either component alone, demonstrating the effectiveness of attention enhancement. And in Table 5, we conduct a comprehensive analysis of the feature aggregation. We explore different sources to provide similarity map. And we find that adjusting both $\mathbb{X}^{penul}$ and $\mathbb{X}^{last}$ is essential, serving as the pre- and post-adjustment for the last layer. Experiments indicate that using layer 9 for pre-adjustment and layer 4 for post-adjustment yields the optimal results, corresponding to (9, 4) in the table, achieving an average 39.8% mIoU, a 6% improvement over the baseline. We hypothesize that deeper mid-layers (e.g., 9), with their broader activations, aggregate richer semantics that help the attention mechanism capture more coherent regions. In contrast, shallower mid-layer (e.g., 4), with more localized activations, can preserve intricate spatial information. And our method is robust to different combinations, with (8, 5) also achieving an mIoU of 39.5. More results are available in the supplementary materials. Furthermore, we compare our approach with DINO, following prior studies [54, 46, 50]. When applied to $\mathbb{X}^{last}$, DINO shows superior performance, showing its fine-grained details; however, our approach achieves comparable results when both pre- and post-adjustments are applied, fully unlocking CLIP’s potential without relying on additional backbones.
Table 5: Ablation study of the feature aggregation.
@@
Table 6: Ablation study of the multi-level fusion strategy.
@@
①baseline 62.1 34.8 37.4 38.0 25.7 39.6
②direct sum 19.5 3.5 2.7 15.4 8.0 9.8
③one pass 63.6 36.1 35.2 39.7 26.1 40.1
two pass 64.6 36.8 37.7 40.1 26.6 41.2
Table 7: Selected layers of the multi-level fusion strategy.
@@
Figure 6: Qualitative Results of Open-Vocabulary Segmentation. We compare our method with CLIP [41], MaskCLIP [74], SCLIP [49] and ClearCLIP [25], all without post-processing. Our SC-CLIP produces much clearer and more accurate segmentation maps.
Multi-level Fusion Strategy

### v3

@@ III-C Self-Adjusting for Semantic Coherence
After resolving the anomaly tokens, the model reduces focus on them in the last layer’s attention. However, a challenge still remains: anomaly tokens have already caused substantial disruption to other normal tokens in the previous layers, diminishing their local awareness.
@@
To alleviate this influence and further enhance feature discriminability and attention correlation, we seek to restore the spatial structure among normal tokens. Some existing methods [34, 68] address this by introducing backbones with strong spatial coherence, such as DINO [9] and SAM [32], to provide fine-grained details. While effective, these approaches rely on additional backbones and incur extra computational costs during training and inference.
@@
Figure 4: Top: Visualization of patch similarities shows that CLIP’s last-layer features perform poorly, but its mid-level features exhibit semantic consistency comparable to DINO. Bottom Left: ROC curve analysis further supports this observation, with our SC-CLIP showing superior semantic coherence. Bottom Right: Detailed ROC analysis of our method.
@@
Figure 5: Illustration of the self-adjusting strategy. (a) We use the similarity map from CLIP’s mid layer to adaptively aggregate deep features by combining semantically similar patches, producing clearer results. The second row provides a detailed process for the selected patch $\bigstar$. (b) We apply the similarity map to enhance attention, broadening and refining the activation regions.
@@
This motivates us to explore whether such semantic coherence can be uncovered within CLIP itself. We begin by visualizing patch similarities, as shown in Figure 4 (top), where we observe that CLIP’s last layer performs poorly as expected. However, its mid-layer exhibits stronger semantic coherence, comparable to that of DINO. To further validate this observation, we quantitatively evaluate all features of CLIP using ROC curve analysis. Specifically, we use 2000 images from the ADE20K validation set [85]. For each image, we extract patch-level features at each layer $l$ and compute their cosine similarity to construct the similarity map $\mathbf{Simi}^{l}\in\mathbb{R}^{N\times N}$, defined as $\mathbf{Simi}^{l}=\frac{\mathbb{X}^{l}\cdot\mathbb{X}^{l}}{\|\mathbb{X}^{l}\|\|\mathbb{X}^{l}\|}$. This similarity map serves as a binary classifier to indicate whether two patches belong to the same category, with patches of the same category labeled as 1 and otherwise 0. Each patch’s category is determined by majority voting on pixel labels in the segmentation map. A higher area under the curve (AUC) in ROC analysis reflects better semantic consistency. As shown in Figure 4 (bottom left), CLIP’s mid-layer features exhibit strong spatial coherence (AUC=0.76), closely matching DINO (AUC=0.77). However, CLIP’s last-layer features perform much worse (AUC=0.66).
@@
The Norm function normalizes the sum to 1. The aggregated features $\hat{\mathbb{X}}^{deep}$ show better results than the original $\mathbb{X}^{deep}$.
@@
Moreover, we find that the self-self attention (e.g., $\mathbf{K}\mathbf{K}^{\top}$ attention) proposed in the previous works [33, 62, 39, 5] exhibits insufficient attention activations. Therefore, we further incorporate the similarity map $\mathbf{Simi}^{mid}$ to augment the attention operation, as detailed in the following formula. As shown in Figure 5 (b), the attention regions now display more extensive and accurate activations.
$\text{attention\_score}=\text{softmax}(\mathbf{K}\mathbf{K}^{\top})+\text{softmax}(\mathbf{Simi}^{mid})$ (5)
@@
Feature aggregation and attention enhancement constitute the self-adjusting strategy, which significantly improves the semantic coherence, increasing the AUC to 0.80, as shown in Figure 4 (bottom right). This demonstrates that the inherent consistency within CLIP’s mid-level features can be effectively leveraged for adjustment.
III-D Two-pass Strategy
@@
Motivated by the effectiveness of multi-level feature aggregation in enriching details for dense prediction tasks [42, 40, 71, 22, 39], we explore the potential of leveraging CLIP’s multi-level features in a training-free manner. A straightforward way is to directly sum features from different layers: $\mathbb{X}^{last}+\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$, where $\mathcal{M}$ denotes a set of mid-level features (e.g., layers 4 to 9). However, CLIP exhibits significant discrepancies across different layers. Specifically, the similarity between $\mathbb{X}^{last}$ and $\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ is merely 0.094, and directly summing them severely disrupts CLIP’s cross-modal alignment capability. To address this issue, we note that different blocks of the ViT-based CLIP encoder capture complementary visual patterns, from local and structural cues in mid-level layers to more global and semantic information in deeper layers. The final visual layer is the only component whose output is directly contrasted with text embeddings during CLIP pre-training, and we therefore assume it effectively acts as an alignment head that maps these visual representations into the joint image–text embedding space. Based on these observations, we derive two key principles for training-free feature fusion: 1 Ensuring the compatibility between $\mathbb{X}^{last}$ and $\sum_{i\in\mathcal{M}}\mathbb{X}^{i}$ is crucial. This can be achieved by leveraging the parameter space of the last layer for alignment. Notably, the similarity between $\mathbb{X}^{last}$ and $\mathbf{L}(\sum_{i\in\mathcal{M}}\mathbb{X}^{i})$ increases to 0.983, where $\mathbf{L}$ denotes the last layer. This indicates that the last layer can effectively realign the aggregated multi-level features to match the original last-layer representation, without any fine-tuning. 2 Preserving the integrity of $\mathbb{X}^{last}$ is critical for maintaining cross-modal alignment, as it directly corresponds to the text embeddings.
@@
Based on these insights, we propose a two-pass strategy, where CLIP’s final layer is explicitly employed for alignment. This strategy involves two forward passes—one with the original $\mathbb{X}^{penul}$ and another with multi-level features, formulated as: $\mathbf{L}(\mathbb{X}^{penul})+\mathbf{L}\left(\sum_{i\in\mathcal{M}}\mathbb{X}^{i}\right).$ Ablation experiments validate that this design enriches the representation with complementary multi-level information while preserving CLIP’s cross-modal alignment capability.
IV Experiment
IV-A Experimental Setup
@@
Datasets and Metric We conduct comprehensive evaluations on eight commonly used benchmark datasets, which are grouped into two categories: 1) with a background class, including PASCAL VOC (VOC21) [20], PASCAL Context (Context) [48], and COCO Object (COCO-Obj) [7]; 2) without a background class, including PASCAL VOC20 (VOC20) [20], Cityscapes (City) [15], PASCAL Context59 (Context59) [48], ADE20K (ADE) [85], and COCOStuff (COCO-Stf) [7]. We evaluate results using the standard mean Intersection-over-Union (mIoU) as the metric.
@@
TABLE I: Performance comparison of our approach with other methods on eight semantic segmentation benchmarks. For a fair comparison, we reproduce all methods following the evaluation protocol in Section IV-A, considering the different settings used by each method. CLIP-DINOiser denotes our reproduced results using OpenAI’s pretrained weights. We report ProxyCLIP* results using its DINO-B/16 variant.
@@
Implementation Details In our experiments, we utilize CLIP [52] with ViT-B/16 and ViT-L/14 architectures. Our code implementation is built on MMSegmentation. After deriving the similarity map $\mathbf{Simi}^{mid}$ from mid-level features, we apply thresholding following previous works [68, 64]. Specifically, values below the threshold $\beta$ are set to zero to strengthen feature correlation, and we set $\beta$ to 0.4. For the evaluation protocol, we adopt the sliding-window inference strategy from SCLIP [62]: input images are resized to have a short side of 336 (560 for Cityscapes [15] due to its higher resolution), and sliding-window inference is conducted with a $224\times 224$ window and a $112\times 112$ stride. No post-processing strategies are applied. Across all datasets, we use the standard ImageNet prompts [52] combined with their category names to construct text descriptions. We keep all hyperparameters consistent across all datasets without separate tuning.
For datasets with the background class (Pascal VOC and COCO-Object), we adhere to the evaluation protocol outlined in previous works [62, 34, 23, 58]. Since the “background” class is overly broad for CLIP and thus challenging to classify, it is represented by a set of stuff categories such as “sky, river, sea, …”, which do not overlap with other target classes. And this pipeline is consistent with prior studies.
TABLE II: Experiments on the MESS benchmark [4].
Method SUIM DeepCrack Kvasir FloodNet Zurich
CLIP-DINOiser [68] 33.5 52.1 48.3 28.2 19.7
ProxyCLIP [34] 35.2 55.0 52.9 31.0 21.1
SC-CLIP (Ours) 36.7 63.1 57.6 33.4 22.0
IV-B Main Results
@@
To ensure a fair comparison, we strictly follow the evaluation protocol specified in the Section IV-A to reproduce all methods. For CLIP-DINOiser [68], we retrain the model using OpenAI’s pretrained weights. Table I presents the comparison of all methods, where our SC-CLIP achieves an average mIoU of 43.9% on CLIP ViT-B/16, setting new state-of-the-art results across eight benchmarks with a notable 9.5% improvement over previous methods. Furthermore, our method demonstrates robustness across different backbones, achieving optimal performance on CLIP ViT-L/14, improving the results by 3.5% average mIoU.
@@
The vanilla CLIP achieves only 14.4% and 6.6% mIoU on ViT-B/16 and ViT-L/14 respectively, demonstrating its limitations in capturing fine-grained spatial details. Our training-free method boosts its performance by threefold on the ViT-B/16 and by 6.8 times on the ViT-L/14.
@@
In contrast to ProxyCLIP [34], which relies on the DINO-B/16 [9] backbone to provide attention weights, our method does not depend on any auxiliary backbone yet achieves superior results. This highlights that the inherent properties of CLIP can be effectively leveraged to calibrate itself, thereby enhancing its feature representation.
The MESS benchmark [4] includes a wide range of domain-specific datasets spanning fields such as earth monitoring, medical sciences, engineering, agriculture, and biology. It serves as a robust evaluation tool for assessing the generalization capability. We conduct experiments on datasets from each domain within the MESS benchmark, and as shown in Table II, our SC-CLIP consistently outperforms CLIP-DINOiser [68] and ProxyCLIP [34], showing superior generalization ability.
IV-C Ablation Study
@@
In this section, we conduct comprehensive ablation experiments to validate the effectiveness of our method. All ablations are performed on the CLIP ViT-B/16 backbone. We use SCLIP [62] as our baseline, which modifies the attention mechanism in the final layer by replacing the original $\mathbf{Q}\mathbf{K}^{\top}$ attention with $\mathbf{Q}\mathbf{Q}^{\top}+\mathbf{K}\mathbf{K}^{\top}$ attention, to enhance correlations. Additionally, we remove the residual connections and feed-forward network (FFN) following [33, 34].
TABLE III: Ablation experiments on the proposed strategies.
@@
+ TwoPass 64.6 36.8 37.7 40.1 26.6 41.2
TABLE IV: Ablation study on resolving the anomaly tokens.
@@
(b) Anomaly detection methods
Isolation Forest 58.9 32.9 34.5 36.0 24.1 37.3
DBSCAN 58.6 33.1 35.1 36.1 24.4 37.5
One-class SVM 59.5 33.2 34.9 36.2 24.3 37.6
LOF 60.3 33.3 34.9 36.5 24.3 37.9
(c) Interpolation methods
Global Mean 56.7 31.9 33.8 34.8 23.3 36.1
CLS token 59.5 32.2 34.0 35.9 23.9 37.1
bilinear 59.9 33.2 34.6 36.3 24.1 37.6
nearest 60.1 33.2 34.7 36.3 24.1 37.7
median 60.3 33.2 34.8 36.5 24.3 37.8
weighted 60.1 33.3 34.8 36.4 24.3 37.8
mean 60.3 33.3 34.9 36.5 24.3 37.9
(d) Neighborhood sizes for interpolation
$\textbf{3}\times\textbf{3}$ 60.3 33.3 34.9 36.5 24.3 37.9
$5\times 5$ 60.1 33.1 34.9 36.4 24.3 37.8
$7\times 7$ 59.8 33.0 34.8 36.3 24.1 37.6
TABLE V: Removing anomaly tokens from different layers.
Methods w/o 1 2 3 4 5 6 7 8 9 10 11 12
Self-self Attn 20.8 0.3 1.8 2.5 0.5 2.6 2.8 2.6 1.6 2.3 13.8 2.0 58.2
Removing Anomaly 58.2 33.5 36.3 40.4 43.5 44.6 45.6 44.7 38.0 8.1 12.4 22.1 60.3
TABLE VI: Ablation study of the attention enhancement.
@@
Only Correlative Self-Self Attention
1 $\mathrm{V}\mathrm{V}^{\top}$ 56.0 30.3 31.3 33.5 22.0 34.6
2 $\mathrm{Q}\mathrm{Q}^{\top}$ 58.4 33.3 35.2 36.6 24.5 37.6
3 $\mathrm{Q}\mathrm{Q}^{\top}$+ $\mathrm{K}\mathrm{K}^{\top}$ 60.3 33.3 34.9 36.5 24.3 37.9
4 $\mathrm{K}\mathrm{K}^{\top}$ 60.6 33.1 34.3 36.6 24.2 37.8
Only $\mathit{Simi}^{mid}$ Attention
5 $\mathrm{Simi}^{mid}$ 60.4 33.8 35.8 37.1 24.8 38.3
Combinations of Both
1 + 5 58.7 32.4 34.0 35.6 23.5 36.8
2 + 5 60.0 33.9 36.1 37.2 24.9 38.4
3 + 5 60.9 33.9 35.9 37.1 24.7 38.5
4 + 5 61.2 34.2 36.2 37.4 25.0 38.8
Analysis of Various Strategies: In Table III, we incrementally incorporate each strategy to highlight its contribution. First, by resolving the anomaly tokens (denoted by AnomRes), our method achieves a 1.2% mIoU improvement, effectively addressing the issues caused by anomaly tokens, as discussed in Section III-B. Next, we apply the self-adjusting strategy proposed in Section III-C, which leverages the spatial consistency within CLIP’s mid-level features to enhance attention correlation (denoted by AttnEnh), leading to a 0.9% improvement. By adaptively aggregating deep features (FeatAgg), we obtain a further 0.8% gain in mIoU. Finally, the two-pass strategy (TwoPass) in Section III-D provides a 1.6% boost in mIoU. Collectively, these strategies contribute to a substantial 12.3% improvement over the baseline. Besides, we conduct a detailed ROC curve analysis as shown in Figure 4 (bottom right). SC-CLIP achieves an AUC of 0.81, surpassing the baseline by 17.4% and significantly enhancing the semantic coherence.
Resolving the Anomaly Tokens: To thoroughly investigate the configurations for resolving anomaly tokens, we conduct a series of ablation studies. These studies examine the impact of varying the number of anomaly tokens, the choice of different anomaly detection methods, various interpolation strategies, and different neighborhood sizes. The results are summarized in Table IV, where the first row reports the baseline results for all four groups of experiments.
@@
(a) The number of anomaly tokens: As we use the LOF algorithm to identify the anomaly tokens, we need to adjust the contamination hyperparameter to control the number of detected anomalies. We explore the optimal number of tokens to resolve, as shown in Table IV (a). As the number increases from 1 to 10, performance gradually improves. However, further increasing this number to 15 yields no additional gains. Therefore, we adopt the removal of 10 tokens (about 5% of the ViT-B/16 token sequence) in our method.
(b) Various anomaly detection methods: We investigate alternative anomaly detection methods to the adopted LOF algorithm, including Isolation Forest, DBSCAN, and One-Class SVM. As shown in Table IV (b), all methods lead to improvements, with LOF algorithm achieving the best results.
(c) Various interpolation methods: We explore various interpolation methods in Table IV (c). After identifying the anomaly tokens, we first try replacing them with the CLS token or the mean of the normal tokens, but these strategies do not lead to performance gains. We then conduct experiments based on the local neighborhood assumption, employing bilinear, nearest-neighbor, median (using the median value of neighboring pixels), weighted (assigning weights based on distance), and mean interpolation. The results show that mean interpolation achieves the best performance, and thus we adopt it as our final choice.
(d) Different neighborhood sizes: As shown in Table IV (d), we compare different neighborhood sizes for mean interpolation. The results show that the $3\times 3$ neighborhood performs better than larger ones. We hypothesize that larger neighborhoods inevitably incorporate more irrelevant regions during the interpolation process, which tends to over-smooth features and blur semantics across object boundaries.
In addition, we also test the removal of anomaly tokens starting from different layers on the VOC21 dataset in Table V. The results show that the modification is only effective when removing from the last layer; otherwise, performance drops significantly. This is because the training-free setting relies on the feature modeling of earlier CLIP layers, and improvements stem from modifying only the final layer. To further support this point, we apply self-self attention [62, 86, 33] (widely validated) to different layers. Results show that it is effective only when applied to the last layer; otherwise, it also drastically degrades performance due to feature space disruption. This suggests that the observed degradation is not caused by LOF’s inability to identify anomaly tokens in other layers.
TABLE VII: Ablation study of the feature aggregation.
@@
Figure 6: Visualization of shallower and deeper mid-layer patch similarity. $\bigstar$ denotes the selected token.
TABLE VIII: Ablation study of the $\beta$ for feature aggregation.
$\beta$ w/o 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Avg 37.4 34.8 37.6 39.3 39.5 38.9 38.2 37.7 37.5 37.4
TABLE IX: Ablation study of the two-pass strategy.
@@
1 baseline 62.1 34.8 37.4 38.0 25.7 39.6
2 direct sum 19.5 3.5 2.7 15.4 8.0 9.8
3 one pass 63.6 36.1 35.2 39.7 26.1 40.1
4 two pass 64.6 36.8 37.7 40.1 26.6 41.2
TABLE X: Selected layers of the two-pass strategy.
@@
Self-adjusting Strategy: To comprehensively evaluate the effectiveness of the self-adjusting strategy, we conduct ablation studies on both attention enhancement and feature aggregation.
Table VI presents the ablation study on attention enhancement, which can be regarded as a more detailed analysis of the “+AttnEnh” entry in Table III. All experiments are conducted with anomaly tokens resolved (i.e., under the “+AnomRes” setting in Table III). We first evaluate different forms of correlative attention ( 1 to 4 in Table VI), and find that these variants achieve comparable performance. We then examine using only $\mathbf{Simi}^{mid}$ ( 5) and its integration with correlative attention. The results show that this integration (Eq. (5)) consistently outperforms either component alone, demonstrating the effectiveness of attention enhancement. Among them, enhancing $\mathbf{K}\mathbf{K}^{\top}$ ( 4+ 5) achieves the best overall performance.
And in Table VII, we conduct a comprehensive analysis of the feature aggregation. We find that adjusting both $\mathbb{X}^{penul}$ and $\mathbb{X}^{last}$ is essential, serving as the pre- and post-adjustment for the last layer. Experiments indicate that using deeper mid-layers (e.g., 9) for pre-adjustment and shallower mid-layer (e.g., 4) for post-adjustment yields the optimal results, corresponding to (9, 4) in the table, achieving an average 39.8% mIoU, a 6% improvement over the baseline. As illustrated in Figure 6, the shallower mid-layers exhibit more localized activations, while deeper mid-layers provide broader activations. Based on this, we hypothesize that deeper layers are well-suited for pre-adjustment due to their broader activations, which help aggregate more regions and enable the attention mechanism to capture more coherent semantics. On the other hand, shallower layers are ideal for post-adjustment, as their localized activations can preserve intricate spatial information. And our method is robust to different combinations, with (8, 3) also achieving an mIoU of 39.7. Furthermore, we compare our approach with DINO, following prior studies [68, 57, 64]. When applied to $\mathbb{X}^{last}$, DINO shows superior performance, showing its fine-grained details; however, our approach achieves comparable results when both pre- and post-adjustments are applied, fully unlocking CLIP’s potential without relying on additional backbones.
In addition, as mentioned in Section IV-A, we apply thresholding to $\mathbf{Simi}^{mid}$, setting values below the threshold $\beta$ to zero. To analyze the effect of the $\beta$ parameter, we experiment with different $\beta$ values using 4th-layer features to adjust $\mathbb{X}^{last}$ (row 8 of Table VII). The results in Table VIII indicate that $\beta$ should not be too large, as this would cause $\mathbf{Simi}$ to activate only for itself. Conversely, if $\beta$ is too small, all tokens would be activated in $\mathbf{Simi}$, leading to feature homogenization. Overall, $\beta$ should be set to a moderate value like 0.4.